Can you please assist? I am not sure I am doing each step correctly. Thanks in advance!!
solve 2(x-4)=2[x-3(x-1)+2]

Answer by ptaylor(2198)

solve 2(x-4)=2[x-3(x-1)+2]
always get rid of the innermost brackets/parens first and we get:
2(x-4)=2[x-3x+3+2] next we'll get rid of the remaining parens/brackets
2x-8=2x-6x+6+4 simplify
2x-8=-4x+10 add 4x and also 8 to both sides
2x+4x-8+8=-4x+4x+10+8 collect like terms
6x=18 divide both sides by 6
x=3
CK
2(3-4)=2[3-3(3-1)+2]
-2=2[3-6+2]
-2=-2
Hope this helps---ptaylor